OpenStudy (anonymous):

WILL FAN AND MEDAL!Retro Rides is a club for owners of vintage cars and motorcycles. Every year the club gets together for a ride. This year, 56 vehicles participated in the ride. The total number of tires of all the vehicles was 156. Assuming each car has 4 tires and each motorcycle has 2 tires, how many each of cars and motorcycles participated in the ride? 22 cars; 34 motorcycles 24 cars; 32 motorcycles 19 cars; 37 motorcycles 26 cars; 30 motorcycles

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

ok You need to form 2 equations

OpenStudy (danjs):

v = vintage cars m = motorcycles

OpenStudy (danjs):

Total vehicals = cars + motorcycles 56 = v + m

OpenStudy (danjs):

Total tires = 4v +2m

OpenStudy (danjs):

solve 56 = v + m 156 = 4v + 2m

OpenStudy (danjs):

22 vehicles and 34 motorcycles

OpenStudy (anonymous):

Mega Movies hosted a film premiere on Friday night. They charged $8 for adults and $6 for children. One hundred thirty-eight adults and children attended, and $1,060 was made in ticket sales. How many children and how many adults went to the film premiere? 19 children; 119 adults 16 children; 122 adults 22 children; 116 adults 26 children; 112 adults

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k

OpenStudy (danjs):

116 adults and 22 children

OpenStudy (danjs):

C

OpenStudy (anonymous):

Sue's Sushi placed two orders with its fish supplier. One order was for 11 pounds of salmon and 6 pounds of tuna; the order totaled $169. The other order was for 8 pounds of salmon and 4 pounds of tuna; this order totaled $120. What is the cost for one pound of salmon and one pound of tuna? $7 salmon; $12 tuna $8 salmon; $11 tuna $13 salmon; $6 tuna $11 salmon; $8 tuna

OpenStudy (danjs):

D

OpenStudy (anonymous):

Which of the following best describes the solution to the system of equations below? 6x + 3y = 5 6x + 3y = 14 The system of equations has exactly one solution where x=5 and y=14 The system of equations has infinitely many solutions. The system of equations has no solution. The system of equations has exactly one solution where x=0 and y=5/3

OpenStudy (danjs):

0 soutions

OpenStudy (anonymous):

Which of the following best describes the solution to the system of equations below? 2x + 6y = 8 6x + 18y = 24 The system of equations has no solution. The system of equations has exactly one solution where x=4 and y=0 The system of equations has exactly one solution where x=1 and y=7/6 The system of equations has infinitely many solutions.

OpenStudy (danjs):

infinite solutions

OpenStudy (danjs):

u good?